OpenStudy (anonymous):

Find the limit when n approaches infinity of cos^2 (n) / 2^n

OpenStudy (anonymous):

I think we have to use Squeeze theorem here... I'm not sure how.

OpenStudy (anonymous):

For all values of \(n\),\[0\le\cos^2n\le1\] So \[\lim_{n\to\infty}0\le\lim_{n\to\infty}\frac{\cos^2n}{2^n}\le\lim_{n\to\infty}2^{-n}\]

OpenStudy (anonymous):

Oh of course, thank you

OpenStudy (anonymous):

yw